OpenStudy (anonymous):

which is irrational? sqrt(32) sqrt(49) 4.05 2/11 Teach me so that I could understand in the future.

OpenStudy (anonymous):

irrational means a number in which the decimals never terminate (can't be formed from a fraction). In this question sqrt 49 simply becomes 7, a number which terminates very nicely. for the sqrt of 32 you can simplify with the following method: break the number into its prime factors, any number you see twice will come outside the sqrt sign as a single value...

OpenStudy (anonymous):

sqrt(32) irrational. irrational numbers are numbers that do not terminate or repeat. repeat

OpenStudy (anonymous):

 \[\sqrt{32}=\sqrt{2*2*2*2*2} \] \[=2*2\sqrt{2}\] \[=4\sqrt2\] sqrt2 is a known irrational number